Locate the absolute maximum and minimum for each of the following functions and justify your responses. show all work!

1) 𝑓(𝑥) = Cube root of x^2 on the interval [−1,1] 

2) 𝑔(𝑥) = 𝑥𝑒^2𝑥 on the interval [−2,0] 


For each of the following functions, respond to the given prompts. Show all work that leads to your responses. Show all work!

 Given 𝑓(𝑥) = 3𝑥 ^3 − 18𝑥^ 2 − 45𝑥 + 10

a. On what interval(s), is 𝑓(𝑥) increasing? Show all work/justify


b. At what value(s) of 𝑥 does 𝑓(𝑥) have a relative minimum? show all work/justify



c. On what interval(s), is 𝑓(𝑥) decreasing and concave up? show all work/justify


1
Expert's answer
2021-02-04T06:19:07-0500

"It\displaystyle
1.\\
(i)\\
f(x) = \sqrt[3]{x^2} = x^{\frac{2}{3}} \\

f'(x) = \frac{2}{3} x^{-\frac{1}{3}} \\


\textsf{At stationary point},\,\, f'(x) = 0 \\

\frac{2}{3} x^{-\frac{1}{3}}= 0,\,\, \frac{2}{3} x^{-\frac{1}{3}} \times x = 0 \times x \\

\frac{2}{3} x^{\frac{2}{3}}= 0

\therefore x = 0


\textsf{Absolute maximum:} \,\,f(1) = f(-1) = \sqrt[3]{1^2} = 1 \\

\textsf{Absolute minimum:} \,\,f(0) = \sqrt[3]{0^2} = 0 \\

(ii)\\ g(x) = xe^{2x} \\

g'(x) = e^{2x} + 2xe^{2x} = e^{2x}(1 + 2x) \\


\textsf{At stationary point,}\,\,g'(x) = 0 \\

e^{2x}(1 + 2x) = 0,\,\, x = -\frac{1}{2} \\


\textsf{Absolute maximum:}\,\, g(0) = 0 \\

\textsf{Absolute minimum:}\,\, g\left(-\frac{1}{2}\right) = -\frac{1}{2e} \\

2\\
(i)\\
f(x) = 3x^3 - 18x^2 - 45x + 10 \\

f'(x) = 9x^2 - 36x - 45 \\

f(x)\,\, \textsf{is increasing if}\,\, f'(x) > 0 \\

9x^2 - 36x - 45 > 0 \\

x^2 - 4x - 5 > 0 \\

(x + 1)(x - 5) > 0 \\

x < -1,\,\, x > 5. \\

(ii)\\
\textsf{At stationary point,}\,\, f'(x) = 0 \\

\therefore x = -1, 5 \\


f''(x) = 18x - 36,\,\, f''(-1) = -54 < 0 \\

f''(5) = 18(5) - 36 = 90 - 36 = 54 > 0\\


\therefore \textsf{By the second derivative test,}\\
\textsf{at}\,\, x = 5\,\, \textsf{there is a minimum.} \\


(iii)\\ 
f(x)\,\, \textsf{is decreasing and concave up}\,\,\textsf{if}\,\, f'(x) < 0. \\

\therefore -1 < x < 5\\

f''(x) = 18x - 36,\\

f(x)\,\, \textsf{is concave up}\,\, \textsf{if}\,\,f"(x) > 0\\\\\n\n18x - 36 > 0 \\\\\n18x > 36\\\\\nx > 2\\\\\n\n\n\\therefore f(x)\\,\\, \\textsf{is concave up}\\,\\, \\textsf{for}\\,\\, x > 2.\\\\\n\n\\textsf{The graph is decreasing at that point}\\\\\n\\textsf{since it lies between}\\,\\, 2 \\,\\, \\textsf{and}\\,\\, 5.\\\\\n\n\n\n\\therefore \\textsf{We can conclude that}\\\\\nf(x)\\,\\, \\textsf{is concave up and decreasing}\\,\\, \\textsf{for}\\,\\, x > 2.\\\\"